OpenStudy (anonymous):

How do I find the equation of a parabola given the vertex and a point on the parabola? One of these problems gives me vertex (0,1) and point (2,2). The book says the answer is x^2 = 4(y-1) (I really don't know where to start because everything I've tried so far hasn't worked.)

jimthompson5910 (jim_thompson5910):

the general form of a parabola is (x-h)^2 = 4p(y-k) where (h,k) is the vertex p is the focal distance (ie the distance from the vertex to the focus along the axis of symmetry)

jimthompson5910 (jim_thompson5910):

the vertex is given to be (0,1) so (h,k) = (0,1) which means h = 0 k = 1

jimthompson5910 (jim_thompson5910):

plug in h = 0 and k = 1 to get (x-h)^2 = 4p(y-k) (x-0)^2 = 4p(y-1) x^2 = 4p(y-1)

jimthompson5910 (jim_thompson5910):

then plug in the other point (x,y) = (2,2), and solve for p, to get x^2 = 4p(y-1) 2^2 = 4p(2-1) 4 = 4p(1) 4 = 4p 4/4 = p 1 = p p = 1

jimthompson5910 (jim_thompson5910):

So, x^2 = 4p(y-1) turns into x^2 = 4*1(y-1) x^2 = 4(y-1)

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (phi):

Jim used the useful form of the equation, but sometimes you see "vertex form" written as \[ y = a(x-h)^2 + k \] (which is the same a Jim's but does not show the "p" ) h and k are the vertex, so you would replace (h,k) with (0,1) \[ y = a(x-0)^2 + 1 \\ y = ax^2 + 1\] you would use the known point (2,2) to find a: \[ 2= a\cdot 2^2 + 1 \\ 1 =4a \\ a= \frac{1}{4} \] and \[ y = \frac{1}{4}x^2 + 1 \] which is "vertex form" of the equation

OpenStudy (anonymous):

Ah, I see. Thank you!